OpenStudy (anonymous):

@johnweldon1993 Sam is driving to Denver. Suppose that the distance to his destination (in miles) is a linear function of his total driving time (in minutes). Sam has 63 miles to his destination after 39 minutes of driving, and he has 51.8 miles to his destination after 53 minutes of driving. How many miles will he have to his destination after 61 minutes of driving?

OpenStudy (johnweldon1993):

A linear function...so he would be traveling at a constant speed...lets see after 39 minutes...he has 63 to go after 53 minutes...he has 51.8 miles to go so in 52 - 39 = 13 minutes he travels 63 - 51.8 = 11.2 miles lets use that to solve how far he drives per minute if he travels 11.2 miles...in 13 minutes....how many miles does he travel in 1 minute?

OpenStudy (anonymous):

0.8615 miles

OpenStudy (johnweldon1993):

right....so now...we know that he only has 51.8 miles left after 53 minutes of driving and now he travels for 61 - 53 = 8 more minutes he has 51.8 - (8 times .8615) miles left to travel do that equation :)

OpenStudy (anonymous):

44.907

OpenStudy (johnweldon1993):

we can round that off to 44.9 ...but yes he will still have 44.9 miles left to go!

OpenStudy (anonymous):

cool thanks :) u r really smart

OpenStudy (johnweldon1993):

no problem! and you're the one doing all the work! :)